Pls help me solve this problem: The units digit of the difference of the squares of two integers is 2.What is the units digits of the sum of these two squares?
Answer by ankor@dixie-net.com(22740)

The units digit of the difference of the squares of two integers is 2.What is the units digits of the sum of these two squares?
:
I am not sure why, but the difference any two squares with units digit of 2,
the sum of these two squares, the units digit is always 0.
:
I can't give a very lucid explanation for it, however
For example:
81 - 49 = 32; 81 + 49 = 130
and
64 - 36 = 22; 64 + 36 = 100
and
16 - 4 = 12; 16 + 4 = 20
and
81 - 9 = 72; 81 + 9 = 90
etc